# Ball Lightning as a profound manifestation of the Dark Matter physics

### v1

@@ 5 Concluding comments and Future Developments
In the present work we focused on manifestations of these annihilation events on possible resolution of the mysterious BL physics. In next Section 5.1 we suggest specific tests which could support or refute the proposed resolution of these BL-related mysteries, while in Sect. 5.2 we overview other manifestations of the same AQN framework at dramatically different scales: from the early Universe to galactic scales, to the solar scale to the Earth scale, where similar mysterious puzzles are known for decades (even centuries), and could be also related to the same antimatter nuggets within the same AQN framework with the same set of parameters.
5.1 Main results and possible future tests of the proposal (17).
We already formulated the basic results demonstrating the consistency of the proposal (17) with observed BL features in Sect. 3.8. We do not need to repeat these results again. Instead, we focus here on possible future tests of the proposal (17). Before we formulate the basic idea for the test we would like to make few comments on observed CR-like events which are very hard to explain within conventional framework and modelling. These events are strongly associated with thunderstorm and lightning events in the area, similar to BL events discussed in this work.
In particular, the Telescope Array (TA) collaboration [41, 42] reported the observations of mysterious bursts when at least three air showers were recorded within 1 ms which cannot occur with conventional high energy CR (it should be days or event months for two or more consecutive energetic CR events hit the same area). These TA mysterious bursts are associated with lightning events in the area. We proposed in [34] that these puzzling events could be related to the AQN-induced events. In fact, in ref. [34] we used the same formula (33) for estimation of the TA mysterious bursts which was used for estimation of the BL frequency of appearance in (36). We explained the rareness of these TA mysterious bursts precisely in the same way as the rareness of BL events.
@@
We mentioned these two studies on unusual CR like events to propose to test the hypothesis (17) as follows. If one can install all sky cameras, similar to the one used in analysis [10], to monitor entire sky in the same area where TA or AUGER detectors (or any other CR labs) are located one can record the BL events with frequency of appearance similar to mysterious bursts recorded by TA collaboration or “Exotic Events” recorded by AUGER collaboration [^7] Similar all sky camera had been used by monitoring entire sky in connection with studies of the meteoroids, and searches for correlations with infrasound signals, see next Sect. 5.2 and ref.[28] with description of another type of mysterious terrestrial events, the so-called sky-quakes, also known for centuries, similar to BL events.. Indeed, in all cases the basic formula describing the rate for all these events is (36). This formula is shown to be consistent with TA mysterious bursts (10 events recorded during 5 years), AUGER “Exotic Events” (23 events recorded during 13 years). The same formula is also consistent with lower bound for BL events (32) as discussed in Sect. 4. Therefore, we predict (within AQN framework) that the rate for BL events in the area will be similar to observed TA mysterious bursts events and AUGER “Exotic Events” because all these events are different manifestations of the same DM physics.
@@
Furthermore, we also predict, within the same AQN framework, that TA mysterious bursts will be correlated in time with observations of the BL in the same area under thunderstorm. The same comment obviously applies to AUGER “Exotic Events”. The recording of such correlations would be very strong argument supporting entire idea about the nature of BL as profound manifestations of the DM physics, as advocated in this work.
@@ 5.2 Other (indirect) evidences for DM in form of the AQN
Another well known puzzle is related to the galaxy formation epoch. The corresponding puzzles commonly are formulated as “Core-Cusp Problem”, “Missing Satellite Problem”, “Too-Big-to-Fail Problem”, to name just a few, see recent reviews [14, 47] for the details [^8] There are many more similar problems and very puzzling observations. We refer to the review papers [14, 47] on this matter. There are also different, but related observations which apparently inconsistent with conventional picture of the structure formation [47].. It has been argued in [48] that the aforementioned discrepancies (and many other related problems referred to in footnote 8) may be alleviated if dark matter is represented in form of the composite, nuclear density objects within AQN frameowrk.
@@
We move from solar scale to terrestrial unusual events. We already mentioned mysterious CR like events in Sect. 5.1. There are many other terrestrial unusual events when conventional picture cannot explain the observed phenomena. In particular, the ANITA observed two anomalous events with non-inverted polarity [56, 57]. Such events correspond to very large inclination angle when “something” crossing the earth before exiting from opposite side of earth at the moment of recording by ANITA. Such events are very hard to explain within conventional physics, but could be explained within AQN framework [58].
Similar to mysterious events mentioned above the AQN can also generate the acoustic and seismic signals, and could be in principle recorded if several dedicated instruments are present in the area. Such events, the so-called sky-quakes have been known (similar to BL) for centuries, without any physical explanations. In fact, a mysterious seismic and atmospheric event has been recorded by a scientific instrument in infrasound frequency band by Elginfield Infrasound Array (ELFO). It has been interpreted in [28] as the AQN-induced phenomenon [^10] One should point out that exceptionally powerful event recorded by ELFO corresponds to the AQN with baryon charge $B\approx 10^{27}$ which is at least $100$ times larger than typical nugget with $B\approx 10^{25}$, see table 1. Such powerful events (similar to the one recorded by ELFO) are expected to be very rare ones. .
@@
We conclude this work with the following final comment. We advocate an idea that the BL known for centuries might be a profound manifestation of the DM physics when DM is made of (anti)quarks and gluons of the Standard Model as reviewed in Sect.2. The model is consistent with all presently available cosmological, astrophysical, satellite and ground-based observations. In fact, it may even shed some light on the long standing puzzles and mysteries as briefly reviewed above. If validated through the proposed tests and experiments, this framework could revolutionize our understanding of DM and its role in the cosmos. The BL as a profound manifestation of the DM on earth (within the AQN framework) could play a key role in the direct study of the DM in scientific labs, in contrast with indirect manifestations of the AQN-induced phenomena reviewed in Sect. 5.2.

### v2

@@ 6 Concluding comments and Future Developments
In the present work we focused on manifestations of these annihilation events on possible resolution of the mysterious BL physics. In next Section 6.1 we suggest specific tests which could support or refute the proposed resolution of these BL-related mysteries. In Section 6.2 we suggest different types of experiments with the instruments which could support or refute our suggestion that some UAP events (such as Pseudo-meteorites, skyquakes, UAV) could be also related to the AQN-induced phenomena as argued in Sect.5. Finally, in Sect. 6.3 we overview other manifestations of the same AQN framework at dramatically different scales: from the early Universe to galactic scales, to the solar scale to the Earth scale, where similar mysterious puzzles are known for decades (even centuries), and could be also related to the same antimatter nuggets (representing the DM objects) within the same AQN framework. We emphasize that the parameters of the model were fixed long ago irrespective to the BL or UAP physics which represents the topic of the present work.
6.1 BL as manifestation of the $AQN_{s}$ events. Possible future tests of the proposal (17).
We already formulated the basic results demonstrating the consistency of the proposal (17) with observed BL features in Sect. 3.8. We do not need to repeat these results again. Instead, we focus here on possible future tests of the proposal (17). Before we formulate the basic idea for the test we would like to make few comments on observed anomalous and mysterious cosmic rays (CR)-like events which are very hard to explain within conventional framework and modelling. These events are strongly associated with thunderstorm and lightning events in the area, similar to BL events discussed in this work.
In particular, the Telescope Array (TA) collaboration [44, 45] reported the observations of mysterious bursts when at least three air showers were recorded within 1 ms which cannot occur with conventional high energy CR (it should be days or event months for two or more consecutive energetic CR events hit the same area). These TA mysterious bursts are associated with lightning events in the area. We proposed in [37] that these puzzling events could be related to the AQN-induced events. In fact, in ref. [37] we used the same formula (33) for estimation of the rate of TA mysterious bursts which was used for estimation of the BL frequency of appearance in (36). We explained the rareness of these TA mysterious bursts precisely in the same way as the rareness of BL events.
@@
We mentioned these two studies on unusual CR like events to propose to test the hypothesis (17) as follows. If one can install all sky cameras, similar to the ones used in analysis [11], to monitor entire sky in the same area where TA or AUGER detectors (or any other CR labs) are located one can record the BL events with frequency of appearance similar to mysterious bursts recorded by TA collaboration or “Exotic Events” recorded by AUGER collaboration [^11] Similar all sky camera had been used by monitoring entire sky in connection with studies of the meteoroids, and searches for correlations with infrasound signals, see Section 5.2.. Indeed, in all cases the basic formula describing the rate for all these events is (36). This formula is shown to be consistent with TA mysterious bursts (10 events recorded during 5 years), AUGER “Exotic Events” (23 events recorded during 13 years). The same formula is also consistent with lower bound for BL events (32) as discussed in Sect. 4. Therefore, we predict (within AQN framework) that the rate for BL events in the area will be similar to observed TA mysterious bursts events and AUGER “Exotic Events” because all these events are different manifestations of the same DM physics.
@@
Furthermore, we also predict, within the same AQN framework, that TA mysterious bursts will be correlated in time with observations of the BL in the same area under thunderstorm. The same comment obviously applies to AUGER “Exotic Events”. The recording of such correlation would be very strong and unambiguous argument supporting entire idea about the nature of BL as profound manifestation of the DM physics, as advocated in this work.
6.2 UAP as manifestation of the AQN events. Possible future tests.
In this subsection we would like to make few comments on possible tests of our interpretation of some UAP events as discussed in Section 5 as the direct manifestation of the AQN events, similar to our interpretation of the BL events as the $AQN_{s}$ events. However, in contrast with proposed tests in previous subsection 6.1 the required instruments to detect the signals from UAP should be very different as they must be sensitive to very different frequency bands.
Therefore, based on the discussions in Sections 5.2 and 5.3 we suggest to test our proposal relating BL, Pseudo-meteorites, skyquakes and UAV phenomena with AQN-induced effects (representing the DM physics in this framework) as follows. First of all, one can search for infrasound acoustic signals, similar to our suggestion formulated in [31] where we proposed to use Distributed Acoustic Sensing (DAS), which is becoming a conventional tool for seismic and other applications. The basic idea of these activities can be explained as follows. It has been known for quite sometime that distributed optical fiber sensors are capable of measuring the signals at thousands of points simultaneously using an unmodified optical fiber as the sensing element. The recent development is that the DAS is capable of measuring strain changes at all points along the optical fiber at acoustic frequencies, which is crucial for our studies of the acoustic waves emitted due to the AQN passage.
Another element which could play a crucial role in identification of the UAV is the radar. We suggest to use two different frequencies simultaneously, one is with larger than plasma frequency, $\omega>\omega_{p}$ and another one is with smaller than plasma frequency $\omega<\omega_{p}$. The radar with $\omega<\omega_{p}$ will detect the image of the ionized cloud produced by the passing AQN, while another radar with $\omega>\omega_{p}$ will be blind to the same area. The recording of the correlation between two radars with different frequencies and DAS should be considered as a very strong and unambiguous argument supporting entire idea about the nature of UAP as profound manifestation of the DM physics in form of the antimatter nuggets powering all these mysterious events. The event rate is determined by the DM flux (33) which is consistent with the BL lower bound as discussed in Section 4, and consistent with observed mysterious and anomalous events (which we believe are consequences of the AQN events) recorded by AUGER and Telescope Array collaborations as reviewed in Section 6.1.
@@ 6.3 Other (indirect) evidences for DM in form of the AQN
Another well known puzzle is related to the galaxy formation epoch. The corresponding puzzles commonly are formulated as “Core-Cusp Problem”, “Missing Satellite Problem”, “Too-Big-to-Fail Problem”, to name just a few, see recent reviews [15, 56] for the details [^12] There are many more similar problems and very puzzling observations. We refer to the review papers [15, 56] on this matter. There are also different, but related observations which apparently inconsistent with conventional picture of the structure formation [56].. It has been argued in [57] that the aforementioned discrepancies (and many other related problems referred to in footnote 12) may be alleviated if dark matter is represented in form of the composite, nuclear density objects within AQN framework.
@@
We move from solar scale to terrestrial unusual events. We already mentioned mysterious CR like events in Sect. 6.1. There are many other terrestrial unusual events when conventional picture cannot explain the observed phenomena. In particular, the ANITA observed two anomalous events with non-inverted polarity [65, 66]. Such events correspond to very large inclination angle when “something” crossing the earth before exiting from opposite side of earth at the moment of recording by ANITA. Such events are very hard to explain within conventional physics, but could be explained within AQN framework [67].
@@
We conclude this work with the following final comment. We advocate an idea that the BL known for centuries might be a profound manifestation of the DM physics when DM is made of (anti)quarks and gluons of the Standard Model as reviewed in Sect.2. We also speculate that other mysterious phenomena known as UAP (it includes pseudo-meteorites, skyquakes and UAV) might also represent different manifestations of the same AQN dark matter physics. The AQN dark matter model is consistent with all presently available cosmological, astrophysical, satellite and ground-based observations. In fact, it may even shed some light on the long standing puzzles and mysteries as briefly reviewed above in Section 6.3. If validated through the proposed tests and experiments, this framework could revolutionize our understanding of DM and its role in the cosmos. The BL and UAP events as profound direct manifestations of the DM on earth (within the AQN framework) could play a key role in the direct study of the DM in scientific labs, in contrast with indirect manifestations of the AQN-induced phenomena reviewed in Sect. 6.3.
@@ Appendix A AQN emission spectrum
Appendix B The BL’s stoppage time from the instant of formation to the moment of observation
The main goal of this Appendix is to provide a quantitative description of the kinematical motion of the $AQN_{s}$ from the instant of formation in thunderclouds at high altitudes to the moment when it starts to behave as a typical BL object emitting the visible frequency light close to the Earth’s surface. We also justify our estimate as given in Sect. 3.6 on typical time scale when BL becomes observable in visible frequency bands.
The process of slowing down of a heavy AQN entering the atmosphere has been discussed previously in the context of the solar corona in [63]. However, it can be directly applied to the $AQN_{s}$ propagating in the Earth’s atmosphere which is the topic of the present work. Therefore, we can generalize formula from [63] to describe the anti-matter $AQN_{s}$ moving with velocity $v_{\rm AQN_{s}}(t)$ in the atmosphere with matter density $\rho_{\rm air}\approx m_{p}n_{\rm air}$:
$M_{\rm AQN_{s}}\frac{dv_{\rm AQN_{s}}(t)}{dt}\approx-\frac{\pi R_{\rm eff}^{2}}{2}\cdot(m_{p}n_{\rm air})\cdot v^{2}_{\rm AQN_{s}}(t),~{}~{}~{}~{}M_{\rm AQN_{s}}\approx m_{p}B_{s}.$ (42)
In what follows we assume that $B_{s}$ varies with time much slower than $v_{\rm AQN_{s}}(t)$, in which case $B_{s}$ can be treated as constant initial baryon charge entering estimate (30) for $L$. In this case the equation (42) can be simplified as follows:
$\frac{dv_{\rm AQN_{s}}(t)}{dt}\approx-\frac{1}{2L}\cdot v^{2}_{\rm AQN_{s}}(t).$ (43)
With all these assumptions and approximations the solution for velocity $v_{\rm AQN_{s}}(t)$ as a function of time $t$ assumes the form
$\left(\frac{1}{v_{\rm AQN_{s}}(t)}-\frac{1}{v_{\rm AQN_{s}}(t_{0})}\right)\approx\frac{(t-t_{0})}{2L},~{}~{}~{}~{}{\rm where}~{}~{}~{}v_{\rm AQN_{s}}(t_{0})\approx v_{\rm AQN}(t_{0})\approx 10^{-3}c$ (44)
Our estimate (44) suggests that a typical stoppage time scale $\tau_{s}$ when velocity $v_{\rm AQN_{s}}(\tau_{s})$ assumes only $\sim 1\%$ of its initial value after spallation can be estimated as follows
$\tau_{s}\approx\frac{2L}{v_{\rm AQN_{s}}(\tau_{s})}\approx\frac{2L}{[0.01\cdot v_{\rm AQN_{s}}(t_{0})]}\approx\frac{6~{}\rm km}{10^{-5}c}\approx 2\rm~{}s,$ (45)
which is precisely the value being used in the main body of the text in Sect. 3.6. The main argument to define stoppage time scale $\tau_{s}$ as the scale when velocity $v_{\rm AQN_{s}}(\tau_{s})$ approximately assumes $\sim 1\%$ fraction from its original value $300~{}\rm km/s$ (when DM object enters the atmosphere) is that the internal temperature of the nugget at this velocity is close to $T\approx 6~{}\rm keV$. Precisely at this typical internal temperatures the BL starts to radiate efficiently in the visible frequency bands as discussed in Sect. 3.3.
Indeed, internal temperature $T$ at the velocity $[0.01\cdot v_{\rm AQN_{s}}]$ assumes the form
$T\sim 20~{}{\rm keV}\cdot\left(\frac{n_{\rm air}}{10^{21}~{}{\rm cm^{-3}}}\right)^{\frac{4}{17}}\cdot\left(\frac{0.01\cdot v_{\rm AQN_{s}}(t_{0})}{v_{\rm AQN}(t_{0})}\right)^{\frac{4}{17}}\approx 6.7~{}\rm keV,$ (46)
and becomes very similar to the temperature (40) we used in our analysis in Sect. 3.3.
The equation (44) allows us to estimate a typical acceleration of the BL after spallation time at $t_{0}$ as follows:
$a_{v_{\rm AQN_{s}}}\equiv\frac{{v_{\rm AQN_{s}}(t)}}{dt}\approx-\frac{2L}{(t-t_{0})^{2}}\approx 1.5\cdot 10^{3}\rm\frac{m}{s^{2}}\cdot\left(\frac{\tau_{s}}{(t-t_{0})}\right)^{2}\gg 10^{2}g.$ (47)
The estimate (47) shows that the acceleration of the $AQN_{s}$ could be enormous and could easily exceed $10^{2}$g as a result of very strong friction as it propagates in atmosphere. In fact it could easily exceed $10^{4}$g at $(t-t_{0})\approx 0.1\tau_{s}\approx 0.2s$ when the velocity of the $AQN_{s}$ already dropped by 90% from its original value to reach $30~{}\rm km/s$. Such enormous acceleration had been mentioned previously in the literature [52] in relation with Unidentified Aerial Phenomena (UAP). We discuss a possible relation between BL and UAP phenomena in the main body of the text in Sect. 5.
Appendix C On ionization of air along the AQN path
The main goal of this appendix is to estimate the plasma frequency $\omega_{p}$ which is the basic element of the discussions of Sect. 5.3. The plasma frequency $\omega_{p}$ characterizes the propagation of photons in the ionized plasma. The $\omega_{p}$ can be thought as an effective mass for photon: only photons (from radar) with the energy larger than this mass can propagate in ionized plasma, while photons with $\omega<\omega_{p}$ will be reflected by ionized plasma back to the radar producing the corresponding image of the ionized cloud. The relevant formula for $\omega_{p}$ reads:
$\omega_{p}^{2}=\frac{4\pi\alpha n_{\rm ion}}{m}.$ (48)
To estimate $\omega_{p}$ we have to estimate the density of the ions $n_{\rm ion}$ which will be produced along the AQN path. The corresponding estimates had been performed in [31] and we use formula (9) from that work
$E_{l}\simeq 10^{4}\cdot\left(\frac{B}{10^{25}}\right)^{2/3}\left(\frac{n_{\rm air}}{10^{21}~{}{\rm cm^{-3}}}\right)\frac{\rm J}{\rm m},$ (49)
where $E_{l}$ has the physical meaning of the energy per unit length injected along the AQN’s path. The energy injection in [31] was estimated at the sea level. For energy injection at high altitudes $z$ one should insert factor $\exp{\left(-{z}/{h}\right)}$ accounting for density decrease with altitude, see (50) below. After that one can estimate the number of emitted photons per unit length along the AQN’s path by dividing $E_{l}$ to the internal temperature of the nugget, i.e. $E_{l}/T$.
The next step is to estimate the photon’s mean free path $L_{\gamma}$ at high altitude $z\approx 8.5\rm~{}km$ where UAV had been observed. It has been also computed in [31] for the sea level altitude and we use formula (12) from that work with adjustment for the altitude dependence:
$L_{\gamma}(z)=\frac{\lambda}{\rho_{\rm air}(z)}\simeq 5{\rm m}\cdot\exp{\left(\frac{z}{h}\right)}\approx 15{\rm m},~{}~{}{\rm where}~{}~{}~{}h\approx 8{\rm~{}km}.$ (50)
The next step is to estimate the density of the ions $n_{\rm ion}$ along the AQN propagation’s path as follows:
$n_{\rm ion}\approx\frac{E_{l}(z)}{[\pi L_{\gamma}^{2}(z)\cdot T]}\approx 2\cdot 10^{9}\rm\frac{ions}{cm^{3}},$ (51)
where we assume that every $T\approx 20\rm~{}keV$ photon (which is our benchmark value) produces a single ion (through the photo-effect as discussed in Sect. 3.3). We consider this estimate as a very conservative assumption as $20\rm~{}keV$ injected energy is capable to produce multiple ions through the chain of processes, such that the $n_{\rm ion}$ could be in fact much higher than estimate (51) suggests.
The final step in our estimate is to substitute (51) into expression (48) for the plasma frequency to arrive
$\omega_{p}=\sqrt{\frac{4\pi\alpha n_{\rm ion}}{m}}\approx 3\cdot 10^{9}\rm s^{-1},$ (52)
which corresponds to the GHz range. This estimate is used in the main body of the text in Sect. 5.3.